OpenStudy (anonymous):

1 1 3 -- + -- = --- 2 x 4 solve for x

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{ 1 }{ x } = \frac{ 3 }{4 }\]

OpenStudy (anonymous):

I'm assuming?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First combine the 1/2 and 3/4 by subtracting 1/2 f rom each side. Because 3/4 - 2/4 = 1/4, 1/x = 1/4

OpenStudy (anonymous):

Then, multiply both sides by x. This gives us: 1 = x/4 4 = x

OpenStudy (anonymous):

The answer is x = 4

OpenStudy (anonymous):

Thank you!!!!!

OpenStudy (anonymous):

Np